# Possibilistic Instrumental Variable Regression

### v1

@@ Abstract
Instrumental variable regression is a common approach for causal inference in the presence of unobserved confounding. However, identifying valid instruments is often difficult in practice. In this paper, we propose a novel method based on possibility theory that performs posterior inference on the treatment effect, conditional on a user-specified set of potential violations of the exogeneity assumption. Our method can provide informative results even when only a single, potentially invalid, instrument is available, offering a natural and principled framework for sensitivity analysis. Simulation experiments and a real-data application indicate strong performance of the proposed approach.
@@ 1 Introduction
We introduce a method designed to accommodate slight violations of the exogeneity assumption. If the instruments are not assumed to be valid a priori, there is no one-to-one correspondence between the (always identifiable) reduced-form parameters and the structural parameters of interest. Based on possibility theory (zadeh_fuzzy_1999; dubois_possibility_2015), our proposed approach can still perform inference on the structural parameters by assigning the “most-possible” reduced-form parameters. We allow the user to specify a set of potential violations and then compute a conditional posterior distribution of the treatment effect given this set. This will generally be uninformative if the set is “too large”, but can be informative if the violations are small. Unlike many existing methods, our method even supports inference with a single invalid instrument.
@@
The main use case we envision for our method is sensitivity analysis. Classical (small_sensitivity_2007; armstrong_sensitivity_2021; cinelli_omitted_2025) as well as Bayesian (or quasi-Bayesian) (conley_plausibly_2012; chernozhukov_plausible_2025) proposals exist in the literature. chib_bayesian_2018 and chernozhukov_plausible_2025 use semi-parametric Bayesian methods to analyse a structural model characterized by a set of moment restrictions, while allowing for the possibility that (some of) these moment conditions do not hold exactly. While chib_bayesian_2018 focus only on misspecification of overidentifying restrictions, chernozhukov_plausible_2025 develop a quasi-Bayesian framework to allow for the possibility that all restrictions are invalid. All approaches that we are aware of in the literature are probabilistic and widen uncertainty intervals to reflect the additional uncertainty about the instruments’ validity. Our contribution is similar, but has the advantage of doing so naturally, as the widened intervals arise directly from epistemic uncertainty about the parameters rather than from ad-hoc adjustments. The literature on partial identification in instrumental variable models is closely related (conley_plausibly_2012; watson_bounding_2024; penn_budgetiv_2025). Our inference is informative only when the invalidity parameter can be constrained to a small set. In that case, all causal effects in the partial identification region have equal posterior possibility. As such, our method emphasises interval estimation rather than point estimation.
@@
The proposed approach is also related to the literature on identification and estimation with invalid instruments. Under the plurality rule, the treatment effect can be identified without knowing in advance which instruments are valid (kang_instrumental_2016). A sufficient condition is that fewer than half of the instruments are invalid. Estimation in this setting typically relies on $\ell_{1}$-penalisation (kang_instrumental_2016; windmeijer_use_2019) or voting/searching strategies (guo_confidence_2018; windmeijer_confidence_2021; guo_causal_2023) to recover the valid instruments. Alternative approaches average across different instrument choices (steiner_bayesian_2025) or assume that the direct effects on the outcome and the treatment are orthogonal (kolesar_identification_2015). In contrast, our approach is not constrained by such restrictive assumptions and can produce informative—though sometimes very diffuse—results even if only a single invalid instrument is available.
@@ 2 Possibility theory
Here, we provide a brief introduction to possibility theory. For more details, we refer to houssineau_parameter_2020; houssineau_robust_2022; hieu_decoupling_2025; martin_possibilistic_2025.
An uncertain variable $\boldsymbol{\theta}$ is a mapping from $\Omega_{\mathrm{u}}\to\Theta$, where $\Omega_{\mathrm{u}}$ is a sample space of deterministic phenomena. We think of $\Omega_{\mathrm{u}}$ as containing a true reference element $\omega_{\mathrm{u}}^{*}$ that is unknown, thus there is no aleatoric uncertainty connected to $\Omega_{\mathrm{u}}$. We characterise the uncertain variable $\boldsymbol{\theta}$ by a possibility function $f_{\boldsymbol{\theta}}:\Theta\to[0,1]$ such that $\sup_{\theta\in\Theta}f_{\boldsymbol{\theta}}(\theta)=1$. This possibility function gives rise to an outer probability measure
@@
where $f_{\boldsymbol{\theta},\boldsymbol{\psi}}$ is a joint possibility function. Marginalising over $\theta$ is done by setting $A=\Theta$ such that the marginal possibility function of $\boldsymbol{\psi}$ is
$f_{\boldsymbol{\psi}}(\psi)=\sup_{\theta\in\Theta}f_{\boldsymbol{\theta},\boldsymbol{\psi}}(\theta,\psi).$
Conditional outer measures can be defined analogously to probability theory as
@@ 3.1 The model
has a unique solution $(\alpha,\beta)$ given $\Gamma$. Without any assumptions on $\alpha$, this is not the case. Typically, one assumes $\alpha=0_{p}$, or that at least the majority of its components are zero (kang_instrumental_2016).
@@ 3.2 Possibilistic inference
More generally, the intersection of the affine subspace $t(\beta)$ and the violation set $A$ defines the partial identification region, in which all values of $\beta$ are equally plausible. To obtain informative results, $A$ must be sufficiently restrictive so that, for most values of $\beta$, the implied $\alpha$ lies outside this region. Figure 1 illustrates this in two dimensions.
[…]
Figure 1: A geometric illustration of our method for $p=2$: The causal effect $\beta$ is partially identified where the affine subspace $t(\beta)=\hat{\gamma}_{1}-\beta\hat{\gamma}_{2}$ intersects the tolerated region AA. At these values of $\beta$, the corresponding $\alpha$ is precisely $t(\beta)$, and therefore, the conditional possibility is $1$. For all other values of $\beta$, the optimal $\alpha$ is the projection onto AA w.r.t. the metric induced by $Z^{\intercal}Z$.
If $A$ is a rectangle, computing the projection is a standard quadratic programming problem. Alternatively, we can also bound a norm of $\alpha$, that is specify the constraint set as $A_{\tau}=\left\{\alpha:\lVert\alpha\rVert\leq\tau\right\}$, where the threshold $\tau$ is the maximum invalidity budget across all instruments (similar to penn_budgetiv_2025). More details are provided in Appendix A.2. In either case, it is important that the choice of $A$ corresponds to the scale of the instrument data. To simplify the interpretation, it may be useful to standardise the instruments.
@@ 3.3 Validification
The sampling properties of our posterior possibility can be improved by using the validification procedure proposed by martin_inferential_2013. Specifically, we transform the posterior possibility defined in Proposition 1 to the validified posterior possibility function
@@
where $F$ is the cdf of a $\chi^{2}$ random variable with $1$ degree of freedom. The Monte Carlo approximation is preferable in small samples, while the $\chi^{2}$ approximation performs well in larger samples. For a comprehensive treatment, we refer to martin_possibilistic_2025. The case with non-vacuous prior information is covered in martin_valid_2023, where the probability measure $P_{\beta}$ has to be replaced by a suitable outer measure.
@@ 4.1 Simulation experiments
Our performance criterion is the empirical coverage of a $95\%$ uncertainty interval for the treatment effect $\beta$. We consider both the $\chi^{2}$ approximation and Monte Carlo (MC) sampling from the validified posterior possibility under different tolerated sets $A$. We compare our empirical coverage to those of naive two-stage least squares (TSLS), plausible generalised method of moments with a Gaussian prior (PGMM-g) (chernozhukov_plausible_2025), and BudgetIV (penn_budgetiv_2025). The results are given in Table 1 and the full details are provided in Appendix B. Code to reproduce our findings is available at https://github.com/gregorsteiner/PossibilisticIV.
@@
In the second experiment, we consider $p=5$ instruments generated from a multivariate standard Gaussian, $s$ of which are invalid. The coefficient $\alpha$ is chosen such that the first $s$ components are $0.1$ and the remaining psp-s components are zero. We set $\gamma_{2}=(1/4,\ldots,1/4)$, which corresponds to a first-stage $R^{2}$ of approximately $1/4$, thus the instrument strength is moderate. The treatment effect is again set to $\beta=1$ and the residuals are generated as above. Now, we also include the confidence interval method (CIIV) (windmeijer_confidence_2021) and gIVBMA (steiner_bayesian_2025).
@@
Table 2 shows the results. For $s=0$, all methods achieve good coverage. As more instruments become invalid, naive approaches, including possibilistic IV with $A=\{0\}$, lose coverage. Our two variants with $A$ containing the true $\alpha$ maintain good coverage even when all instruments are invalid. However, when $\alpha$ lies on a corner of the hypercube (e.g., $A=[0.0,0.2]^{p}$ with $s=0$ or $A=[-0.1,0.1]^{p}$ with $s=5$), the intervals can be slightly overoptimistic, particularly for the $\chi^{2}$ approximation. BudgetIV is the only other method that can maintain coverage above the nominal level for $s=5$.
Our approach widens the confidence sets appropriately when the instruments are not assumed to be valid a priori. This ensures valid inference even if no valid instruments exist. These sets may include a non-unique mode, reflecting partial identification. However, as shown in the experiments above, choosing $A$ too wide can make the confidence sets overly conservative and less useful in practice. BudgetIV can similarly maintain good coverage, as it also relies on partial identification, yet it may not return a plausible set for certain hyperparameter specifications. In contrast, our method always yields a posterior possibility function, even if it is sometimes very uninformative.
4.2 A real data example: The effect of institutions on economic growth
We illustrate our method with an empirical application estimating the effect of institutions on economic output. The analysis uses the dataset of 64 countries originally compiled by acemoglu_colonial_2001 and reexamined by chernozhukov_plausible_2025 to demonstrate their quasi-Bayesian approach that allows for small violations of instrument exogeneity.
The outcome is log GDP per capita in 1995, with the main predictor being protection against expropriation, a proxy for institutional quality. A central challenge is endogeneity: institutions may raise income, but income may also improve institutions. To address this, acemoglu_colonial_2001 use the mortality of early European settlers as an instrument, arguing that long-term settlement incentives shaped institutional quality. The instrument’s validity rests on the assumption that settler mortality is exogenous (conditional on covariates). Following chernozhukov_plausible_2025, a direct effect of settler mortality should not be stronger than $0.1$ in absolute terms, if any, justifying A=[0.1,0.1]A=[-0.1,0.1] as a reasonable violation set. We consider the specification with an intercept and (normalised) distance from the equator as exogenous control variables and (log) settler mortality as the sole instrument. We project out the covariates and run our analysis on the residuals resulting from that projection.
@@ 4.2 A real data example: The effect of institutions on economic growth
Figure 2 shows validified posterior possibility functions for the valid case and allowing some invalidity. The curve corresponding to A=[0.1,0.1]A=[-0.1,0.1] is wider and lacks a unique mode, reflecting the partial identification of $\beta$. Both indicate a significantly positive effect with relatively heavy right tails, consistent with chernozhukov_plausible_2025. For comparison, we also include the violation set A=[0.4,0.4]A=[-0.4,0.4], which is wide enough for the significant effect to disappear. For $A=\{0\}$, the $\chi^{2}$ and MC approximations look essentially identical. For larger $A$, however, the latter display steeper decay away from the partial identification region. Table 3 shows that our uncertainty intervals are narrower than those of chernozhukov_plausible_2025, especially on the right tail. Relaxing the exogeneity assumption and allowing for $\alpha\in[-0.1,0.1]$ does not qualitatively change the conclusion that good institutions promote economic output.
Table 3: The effect of institutions on economic growth: $95\%$ uncertainty intervals for possibilistic IV, TSLS, and PGMM (taken from chernozhukov_plausible_2025). The intervals for our approach are based on the $\chi^{2}$ approximation. PGMM-u, PGMM-g and PGMM(d)-g refer to PGMM with, respectively, uniform prior, baseline Gaussian prior and diffuse Gaussian prior.
Method $95\%$ Interval
Possibilistic IV ($A=\{0\}$) $[0.69,1.62]$
Possibilistic IV (A=[0.1,0.1]A=[-0.1,0.1]) $[0.53,1.96]$
TSLS $[0.56,1.38]$
PGMM-u $[0.58,3.65]$
PGMM-g $[0.49,3.79]$
PGMM(d)-g $[0.22,3.81]$
@@
To further interpret the obtained posterior possibility, we can view it as an upper bound on a precise (subjective) probability measure. This allows to deduce a corresponding lower bound, and therefore provides a probability interval for the event of interest. Table 4 displays such intervals for the hypothesis $\beta>0$, given by the pair $\left(1-\sup_{\beta\leq 0}\pi_{w}(\beta\mid A),\sup_{\beta>0}\pi_{w}(\beta\mid A)\right)$, under different choices of the violation set $A$. It takes $\alpha$ being close to $0.4$ in absolute terms to materially change the inference, that is, the lower probability drops below the nominal level. Thus, the qualitative effect of institutions on economic output is very robust to reasonable violations of the exogeneity assumptions.
Table 4: The effect of institutions on economic growth: Lower and upper probabilities for the hypothesis $\beta>0$ under the constraint $\alpha\in A$ (based on the MC approximation with $M=10,000$).
A Lower Upper
$\{0\}$ 1 1
[0.1,0.1][-0.1,0.1] 1 1
[0.2,0.2][-0.2,0.2] 1 1
[0.3,0.3][-0.3,0.3] 1 1
[0.4,0.4][-0.4,0.4] 0.927 1
[0.5,0.5][-0.5,0.5] 0.603 1

### v2

@@ Abstract
Instrumental variable regression is a common approach for causal inference in the presence of unobserved confounding. However, identifying valid instruments is often difficult in practice. In this paper, we propose a novel method based on possibility theory that performs posterior inference on the treatment effect, conditional on a user-specified set of potential violations of the exogeneity assumption. Our method can provide valid results even when only a single, potentially invalid, instrument is available. A theoretical coverage guarantee for our uncertainty intervals holds if the violation set contains the true value. Simulation experiments and real-data applications indicate strong performance of the proposed approach.
@@ 1 Introduction
We introduce a method designed to accommodate slight violations of the exogeneity assumption. If the instruments are not assumed to be valid a priori, there is no one-to-one correspondence between the (always identifiable) reduced-form parameters and the structural parameters of interest. Based on possibility theory (Zadeh,, 1999; Dubois and Prade,, 2015), our proposed approach can still perform inference on the structural parameters by assigning the “most-possible” reduced-form parameters. We allow the user to specify a set of potential violations and then compute a conditional posterior distribution of the treatment effect given this set. This will generally be uninformative if the set is “too large”, but can be informative if the violations are small. Unlike many existing methods, our method even supports inference with a single invalid instrument. Based on the possibilistic inferential model construction (Martin and Liu,, 2013; Martin,, 2025), we prove a frequentist coverage guarantee for our uncertainty intervals that holds if the violation set is chosen correctly.
@@
An important use case we envision for our method is sensitivity analysis. Classical (Small,, 2007; Armstrong and Kolesár,, 2021; Cinelli and Hazlett,, 2025) as well as Bayesian (or quasi-Bayesian) (Conley et al.,, 2012; Chernozhukov et al.,, 2025) proposals exist in the literature. Chib et al., (2018) and Chernozhukov et al., (2025) use semi-parametric Bayesian methods to analyse a structural model characterized by a set of moment restrictions, while allowing for the possibility that (some of) these moment conditions do not hold exactly. While Chib et al., (2018) focus only on misspecification of overidentifying restrictions, Chernozhukov et al., (2025) develop a quasi-Bayesian framework to allow for the possibility that all restrictions are invalid. All approaches that we are aware of in the literature are probabilistic and widen uncertainty intervals to reflect the additional uncertainty about the instruments’ validity. Our contribution is similar, but has the advantage of doing so naturally, as the widened intervals arise directly from epistemic uncertainty about the parameters rather than from ad-hoc adjustments. The literature on partial identification in instrumental variable models is closely related (Conley et al.,, 2012; Watson et al.,, 2024; Penn et al.,, 2025). Our inference is informative only when the invalidity parameter can be constrained to a small set. In that case, all causal effects in the partial identification region have equal posterior possibility. As such, our method emphasises interval estimation rather than point estimation.
@@
The proposed approach is related to the literature on identification and estimation with invalid instruments. Under the plurality rule, the treatment effect can be identified without knowing in advance which instruments are valid (Kang et al.,, 2016). A sufficient condition is that fewer than half of the instruments are invalid. Estimation in this setting typically relies on $\ell_{1}$-penalisation (Kang et al.,, 2016; Windmeijer et al.,, 2019) or voting/searching strategies (Guo et al.,, 2018; Windmeijer et al.,, 2021; Guo,, 2023) to recover the valid instruments. Alternative approaches average across different instrument choices (Steiner and Steel,, 2025) or assume that the direct effects on the outcome and the treatment are orthogonal (Kolesár et al.,, 2015). In contrast, our approach is not constrained by such restrictive assumptions and can produce informative—though sometimes very diffuse—results even if only a single invalid instrument is available.
@@ 2 Possibility theory
Here, we provide a brief introduction to possibility theory. For more details, we refer to Houssineau, (2020); Houssineau and Nott, (2022); Hieu et al., (2025); Martin, (2025).
An uncertain variable $\boldsymbol{\theta}$ is a mapping from $\Omega_{\mathrm{u}}\to\Theta$, where $\Omega_{\mathrm{u}}$ is a sample space of deterministic phenomena. We think of $\Omega_{\mathrm{u}}$ as containing a true (but unknown) reference element $\omega_{\mathrm{u}}^{*}$, thus there is no aleatoric uncertainty connected to $\Omega_{\mathrm{u}}$. We characterise the uncertain variable $\boldsymbol{\theta}$ by a possibility function $f_{\boldsymbol{\theta}}:\Theta\to[0,1]$ with $\sup_{\theta\in\Theta}f_{\boldsymbol{\theta}}(\theta)=1$. This possibility function gives rise to an outer probability measure
@@
where $f_{\boldsymbol{\theta},\boldsymbol{\psi}}$ is a joint possibility function. Marginalising over $\theta$ is done by setting $A=\Theta$ such that the marginal possibility function of $\boldsymbol{\psi}$ is $f_{\boldsymbol{\psi}}(\psi)=\sup_{\theta\in\Theta}f_{\boldsymbol{\theta},\boldsymbol{\psi}}(\theta,\psi)$. Conditional outer measures can be defined analogously to probability theory as
@@ 3.1 The model
has a unique solution $(\alpha,\beta)$ given $\Gamma$. Without any assumptions on $\alpha$, this is not the case. Typically, one assumes $\alpha=0_{p}$, or that at least the majority of its components are zero (Kang et al.,, 2016).
@@ 3.2 Possibilistic inference
More generally, the intersection of the affine subspace $t(\beta)$ and the violation set $A$ defines the partial identification region, in which all values of $\beta$ are equally plausible. To obtain informative results, $A$ must be sufficiently restrictive so that, for most values of $\beta$, the implied $\alpha$ lies outside this region.
If $A$ is a rectangle, computing the projection is a standard quadratic programming problem. Alternatively, we can also bound a norm of $\alpha$, that is specify the constraint set as $A_{\tau}=\left\{\alpha:\lVert\alpha\rVert\leq\tau\right\}$, where the threshold $\tau$ is the maximum invalidity budget across all instruments (similar to Penn et al.,, 2025). More details are provided in Appendix A.2. In either case, it is important that the choice of $A$ corresponds to the scale of the instrument data. To simplify the interpretation, it may be useful to standardise the instruments.
@@ 3.3 Validification
The sampling properties of our posterior possibility can be improved by using the validification procedure proposed by Martin and Liu, (2013). Specifically, we transform the posterior possibility defined in Proposition 1 to the validified posterior possibility function
@@
where $F$ is the cdf of a $\chi^{2}$ random variable with $1$ degree of freedom. This approximation is based on asymptotic Gaussianity and is less appropriate whenever the parameter is not point-identified, i.e., when $A$ is not a singleton. For a comprehensive treatment, we refer to Martin, (2025). The case with non-vacuous prior information is covered in Martin, (2023), where the probability measure $P_{\beta}$ has to be replaced by a suitable outer measure.
@@ 4.1 Simulation experiments
Our performance criterion is the empirical coverage of a $95\%$ uncertainty interval for the treatment effect $\beta$. We consider both the $\chi^{2}$ approximation and Monte Carlo (MC) sampling from the validified posterior possibility under different tolerated sets $A$. We compare our empirical coverage to those of naive two-stage least squares (TSLS), plausible generalised method of moments with a Gaussian prior (PGMM-g) (Chernozhukov et al.,, 2025), and BudgetIV (Penn et al.,, 2025). The results are given in Table 1 and the full details are provided in Appendix B. Code to reproduce our findings is available at https://github.com/gregorsteiner/PossibilisticIV.
@@
In the second experiment, we consider $p=5$ instruments generated from a multivariate standard Gaussian, $s$ of which are invalid. The coefficient $\alpha$ is chosen such that the first $s$ components are $0.1$ and the remaining psp-s components are zero. We set $\gamma_{2}=(1/4,\ldots,1/4)$, which corresponds to a first-stage $R^{2}$ of approximately $1/4$, thus the instrument strength is moderate. The treatment effect is again set to $\beta=1$ and the residuals are generated as above. Now, we also include the confidence interval method (CIIV) (Windmeijer et al.,, 2021) and gIVBMA (Steiner and Steel,, 2025).
@@
Table 2 shows the results. For $s=0$, all methods achieve good coverage. As more instruments become invalid, naive approaches, including possibilistic IV with $A=\{0\}$, lose coverage. Our two variants with $A$ containing the true $\alpha$ maintain good coverage even when all instruments are invalid. However, when $\alpha$ lies on a corner of the hypercube (e.g., $A=[0.0,0.2]^{p}$ with $s=0$ or $A=[-0.1,0.1]^{p}$ with $s=5$), the intervals can be slightly too short, particularly for the $\chi^{2}$ approximation. BudgetIV is the only other method that can maintain coverage above the nominal level for $s=5$.
Our approach widens the confidence sets appropriately when the instruments are not assumed to be valid a priori. This ensures valid inference even if no valid instruments exist. These sets may include a non-unique mode, reflecting partial identification. However, as shown in the experiments above, choosing $A$ too wide can make the confidence sets overly conservative and less useful in practice. BudgetIV can similarly maintain good coverage, as it also relies on partial identification, yet it may not return a plausible set for certain hyperparameter specifications (see Appendix B). In contrast, our method always yields a posterior possibility function, even if it is sometimes very uninformative.
4.2 The effect of institutions on economic growth
We illustrate our method with an empirical application estimating the effect of institutions on economic output. The analysis uses the dataset of 64 countries originally compiled by Acemoglu et al., (2001) and reexamined by Chernozhukov et al., (2025) to demonstrate their quasi-Bayesian approach that allows for small violations of instrument exogeneity.
The outcome is log GDP per capita in 1995, with the main predictor being protection against expropriation, a proxy for institutional quality. A central challenge is endogeneity: institutions may affect income, but income may also influence institutions. To address this, Acemoglu et al., (2001) use the mortality of early European settlers as an instrument, arguing that long-term settlement incentives shaped institutional quality. The instrument’s validity rests on the assumption that settler mortality is exogenous (conditional on covariates). Following Chernozhukov et al., (2025), a direct effect of settler mortality should not be stronger than $0.1$ in absolute terms, if any, justifying A=[0.1,0.1]A=[-0.1,0.1] as a reasonable violation set. We consider the specification with an intercept and (normalised) distance from the equator as exogenous control variables and (log) settler mortality as the sole instrument. We project out the covariates and run our analysis on the residuals resulting from that projection.
@@ 4.2 The effect of institutions on economic growth
Figure 1 shows validified posterior possibility functions for the valid case and allowing some invalidity. The curve corresponding to A=[0.1,0.1]A=[-0.1,0.1] is wider and lacks a unique mode, reflecting the partial identification of $\beta$. Both indicate a significantly positive effect with heavier right tails, consistent with Chernozhukov et al., (2025). For comparison, we also include the violation set A=[0.4,0.4]A=[-0.4,0.4], which is wide enough for the significant effect to disappear. Table C.1 in Appendix C shows that our uncertainty intervals are narrower than those of Chernozhukov et al., (2025), especially on the right tail. Relaxing the exogeneity assumption and allowing for $\alpha\in[-0.1,0.1]$ does not qualitatively change the conclusion that good institutions promote economic output.
For $A=\{0\}$, the $\chi^{2}$ and MC approximations look essentially identical. For larger $A$, however, the latter display steeper decay away from the partial identification region. This is expected as the Gaussian approximation cannot correctly capture the tail behaviour when $\beta$ is not point-identified. In these cases, we observe that the $\chi^{2}$ approximation leads to more conservative (but still valid) inference. This suggests a practical strategy can be fitting the $\chi^{2}$ approximation first, and then only fitting the computationally expensive MC approximation if additional tightness is required.
@@
To further interpret the obtained posterior possibility, we can view it as an upper bound on a precise (subjective) probability measure. This allows to deduce a corresponding lower bound, and therefore provides a probability interval for the event of interest. Table C.2 in the Appendix displays such intervals for the hypothesis $\beta>0$, given by the pair $\left(1-\sup_{\beta\leq 0}\pi_{w}(\beta\mid A),\sup_{\beta>0}\pi_{w}(\beta\mid A)\right)$, under different choices of the violation set $A$. It takes $\alpha$ being close to $0.4$ in absolute terms to materially change the inference, that is, the lower probability drops below the nominal level. Thus, the qualitative effect of institutions on economic output is very robust to reasonable violations of the exogeneity assumptions.
4.3 The returns to schooling
We revisit the relationship between education and wages using the dataset from Card, (1995). Education choices are not random, but are influenced by unobserved traits that also affect earnings, so the observed link between schooling and wages may not reflect the true causal effect. Card, (1995) adresses this endogeneity by using college proximity as an instrumental variable. The outcome is (the logarithm of) hourly wages and the treatment is years of schooling received, which we instrument with proximity to a four-year college. We partial out the available covariates, including experience, experience squared, family background, marital status, race, region, and parents’ education. The parental education variables are missing for a substantial proportion of the sample. Following Card, (1995), we impute them using the mean and include an indicator for missingness.
We also expect the direct effect of proximity to a college on hourly wages to be positive, if it exists. A magnitude of the effect in between zero and $4\%$ seems reasonable to us. Thus, we consider the violation sets $A=[0,0.02]$ and $A=[0,0.04]$. Figure 2 presents the results. For $A=\{0\}$, the posterior mode is just below $0.15$, comparable to the estimates in Card, (1995), and the effect is significant, since the uncertainty intervals do not include zero. In the other two cases, the left tail is shifted to the left and the effect is no longer significant, since the uncertainty intervals now include zero. Thus, positive returns to education are rather sensitive to violations of the exogeneity assumption of the instrument “college proximity”. Finally, under the MC approximation, the right tail is thinner, assigning lower possibility to very high returns to schooling.
Figure 2: The returns to schooling: Validified posterior possibility functions under a perfectly valid instrument ($\alpha=0$) and allowing for potential violations $A=[0,0.02]$ and $A=[0,0.04]$. The solid line is based on the $\chi^{2}$ approximation, while the dashed line displays the Monte Carlo approximation. The dashed grey line indicates the 0.05 level, such that the 95% uncertainty interval for $\beta$ includes all values where the posterior possibility function exceeds this threshold.
As in the previous example, we can compute lower and upper probabilities for the hypothesis that the returns to schooling are positive, see Table C.3 in the Appendix. We see that the positive treatment effect is robust to a direct effect of college proximity on wages of $1\%$, whereas larger violations lead the lower probability to drop significantly. For a direct effect of $4\%$, the lower probability of positive returns to schooling is only approximately $0.52$.